OpenStudy (anonymous):

Pre-Calc : I've been looking at my test that I got back today, and I'm not sure how i got this one wrong. The function is g(x) = abs(x) if x less than or equal to 0, and x^2 if x is greater than 0. The problems I got wrong were g(-2.3) and g(4.5), and I'd appreciate if someone answered these two with explanation :)

OpenStudy (anonymous):

Well, we know that g(x) is a peace-wise function so let's state that here.. g(x) = abs(x) | x <= 0 g(x) = x^2 | x > 0 so part a) We know that x = -2.3, therefore we use the first case where x <= 0 and say g(-2.3) = abs(-2.3) = 2.3 and part b) We know that x = 4.5, and thus use the second case where x > 0 and say g(4.5) = (4.5)^2 = 20.25 Let me know if you don't understand.

OpenStudy (anonymous):

Ohhhh Thank you! I plugged in -2.3 to x^2 and the 4.5 to abs(x).. Stupid mistake x_x.. Thank you for the explanation! :)